OpenStudy (pinterestaddict15):

Will give medal. The legs of an isosceles triangle have lengths x + 1 and -x + 7 . The base has length 3x - 3. What is the length of the base? A. 4 B. 6 C. 3 D. Cannot be determined.

sammixboo (sammixboo):

This one @pinterestaddict15 ?

OpenStudy (pinterestaddict15):

Yes, I put C and got it wrong.

sammixboo (sammixboo):

I think it is D, because you do not have enough information. Can you try that? I am 90% sure it is D

OpenStudy (pinterestaddict15):

Okay I'll try that, thanks, I suck at math.

Directrix (directrix):

@pinterestaddict15 By definition, the legs of an isosceles triangle are congruent. That means that the legs have the same length. Therefore, x + 1 and -x + 7. Solving, indeed you do get x = 3 as you cranked out above. Then, you posted "Yes, I put C and got it wrong."

Directrix (directrix):

The question is to find the length of the base, NOT the value of x. So, if x = 3, then the base 3x - 3 = 3*3 - 3 = ? I see it as one of the options. What is it @pinterestaddict15